# USPilot: An Embodied Robotic Assistant Ultrasound System with Large Language Model Enhanced Graph Planner

### v1

@@ Abstract
In the era of Large Language Models (LLMs), embodied artificial intelligence presents transformative opportunities for robotic manipulation tasks. Ultrasound imaging, a widely used and cost-effective medical diagnostic procedure, faces challenges due to the global shortage of professional sonographers. To address this issue, we propose USPilot, an embodied robotic assistant ultrasound system powered by an LLM-based framework to enable autonomous ultrasound acquisition. USPilot is designed to function as a virtual sonographer, capable of responding to patients’ ultrasound-related queries and performing ultrasound scans based on user intent. By fine-tuning the LLM, USPilot demonstrates a deep understanding of ultrasound-specific questions and tasks. Furthermore, USPilot incorporates an LLM-enhanced Graph Neural Network (GNN) to manage ultrasound robotic APIs and serve as a task planner. Experimental results show that the LLM-enhanced GNN achieves unprecedented accuracy in task planning on public datasets. Additionally, the system demonstrates significant potential in autonomously understanding and executing ultrasound procedures. These advancements bring us closer to achieving autonomous and potentially unmanned robotic ultrasound systems, addressing critical resource gaps in medical imaging.
@@ I Introduction
Ultrasound is widely used in medical imaging. It provides real-time images of internal organs, blood vessels, and tissues without exposing patients to ionizing radiation. This technique is particularly useful for diagnosing conditions such as pregnancy, heart disease, and tumors. However, hospital sonographers face significant challenges due to the increasing demand for ultrasound services and the rapid pace of technological advancements. There are over 64,751 sonographers employed in the United States, but the number will need to increase by 27,600 according to a prediction by the Bureau of Labor Statistics[1]. In addition, sonographers commonly experience physical strain from their work. Repetitive motions often lead to joint discomfort, particularly in the wrists, upper back, and shoulders[2]. Research on diagnostic medical sonographers indicates that a significant majority—up to 91%—experience discomfort or pain while conducting ultrasound examinations[3]. To mitigate this issue, it is imperative to consider the implementation of automation, particularly in light of the global shortage of ultrasound imaging professionals and the increasing demand for ultrasound-based diagnostic imaging.
The introduction of automated diagnostic and intraoperative ultrasound systems, which are either guided or assisted by robotic technology, represents a promising solution to the challenges previously outlined. By diminishing the workload of ultrasound technicians, these automated systems pave the way for more sophisticated image acquisition processes characterized by digital and numerical consistency. There are currently several robotic assistant US scanning systems for specific tasks [4] such as reinforcement learning[5, 6], imitation learning [7, 8] and predefined tasks[9]. However, these systems still require extensive domain knowledge or reward engineering, requiring doctors or engineers to select the specific functions individually. The human-robot interaction methods in these systems may also increase surgeons’ workload or distract them from their current tasks, which do not align with their operating habits.
@@
With the rise of large language models (LLMs), a subsequent surge in embodied artificial intelligence has been observed. LLMs have learned rich semantic knowledge of the world, so they can be naturally used for planning without complex definitions or additional training. This ability of LLMs to understand and generate human-like text based on vast amounts of training data allows them to be seamlessly integrated into various applications where decision-making and planning are crucial. Some research, such as Octopus [10] and TAPA [11], illustrated the ability of task reasoning and planning for daily instructions in simulated environments. Some research has trained on robot manipulation data and acting the robot arm with the pre-defined vision feedback function [12, 13, 14]. [15] demonstrating its ability to perform common surgical sub-tasks through human language instruction. However, these methods achieve low-level robot control by using LLMs to code under the guidance of prompt engineering, which will face the illusion of the LLMs. The illusion of LLMs in low-level robot control coding may cause safety issues, especially in medical robotics applications.
In this paper, we present USPilot, a comprehensive framework designed for autonomous ultrasound acquisition. The framework incorporates a semantic router powered by Large Language Models (LLMs) that interpret user queries related to ultrasound procedures and translate natural language into instructions for robotic ultrasound system operation. A key innovation of our framework is the implementation of an LLM-enhanced Graph Neural Network (GNN) for robotic tool selection and planning. By leveraging adapter-based ultrasound knowledge integration, USPilot demonstrates successful autonomous ultrasound image acquisition on a physical robotic platform, marking a significant step toward fully autonomous robotic ultrasound systems. This advancement suggests promising possibilities for unmanned medical imaging applications.
II Related works
@@
LLMs as Planners For robots, language serves as a natural medium to guide high-level control. Many methods have attempted to apply the prior knowledge of LLMs and Vision Language Models (VLMs) to robot tasks. These models can assist in breaking down long-horizon tasks[16] and responding in interactive environments. Techniques like Chain-of-Thought[17] and few-shot prompting are useful for ensuring that each step of a plan is well-reasoned and atomic. However, LLM illusions and associations can still introduce instability and errors, which is unacceptable in ultrasound tasks. To address these issues, researchers have designed prompts or trained LLMs to adapt to frameworks like linear temporal logic[18], code generation[19, 20, 16], and behavior trees[21]. These constraints help preserve associations while ensuring stability. In our approach, we use a text-attributed graph (TaG) to constrain the task space for LLM planning.
TAG with LLMs In TaGs, textual descriptions of entities act as vertexes. At the same time, the edges represent relationships between neighbors, which, in our case, correspond to robot tasks in the ultrasound scan and their execution logic. To leverage the power of LLMs on the text in graphs, existing methods assign LLMs three roles: aligner, predictor, and enhancer. Researchers[22, 23] align the output from Graph Neural Networks (GNN) with LLMs to create a shared latent space for knowledge transformation. Similar results can be achieved using a projector applied to the GNN encoder, like GraphGPT[24, 25], which helps make the graph embedding more understandable for the LLM predictor. However, aligner LLMs can introduce illusions into graphs, and LLMs as predictors require high-quality labeled datasets to ensure accurate projection. Since LLMs can significantly enhance or refine textual features, they are used in[26, 27] to improve text features for GNNs. In our ultrasound robot task, this structure enriches the original task information while limiting the LLM cognition domain.
III Methodology
III-A Problem Formulation
@@
Robotic ultrasound manipulation can be represented as a task planning problem on a graph, where each step in the system is highly dependent on its preceding steps. The graph is defined as a TaG with the structure $G=(V,E,T)$. The vertex $v\in V$ represents a task in the ultrasound scan, which is defined as an application programming interface (API) in the robotic system. The API description text $t\in T$ provides a linguistic description of the function of each API. The edge $(x,y)\in E$ indicates the dependency between two APIs. The graph-based task planning aims to identify an optimal subgraph $G^{*}$ that can complete the given task instruction $I$.
@@ III-B Structure Overview
The overview of USPilot is depicted in Fig 2. The semantic router is a dynamic routing mechanism connected to the LLM transformer structure. It enables user intent recognition, which helps the system load the appropriate knowledge or graphs. Related subject knowledge is trained into the LLM through an adapter structure, allowing the model to better respond to user intent by adjusting the gating factor of the adapter layers.
@@
If the user intent involves executable tasks, the router switches to the Large Language Model Enhanced Graph (LLMEG). The LLMEG comprises an LLM-enhanced Graph Neural Network (GNN) head and a subgraph generator. The LLM-enhanced GNN head selects APIs by sending LLM-generated word embeddings and GNN-generated graph embeddings to the decoder. The LLM-based subgraph generator defines the path for API execution. When the user intent involves a general medical question, the router switches the inference path to a pre-trained ultrasound-related question-answer adapter to fulfill the medical inquiry requirements. The semantic router, adapters, and LLMEG are all hot-plugging in our USPilot, which makes it suitable for separate training, upgrading, and adapting.
III-C Large Language Model Enhanced Graph
@@
LLMs possess a remarkable ability to interpret user requests and decompose complex commands into simpler subtasks. This task decomposition capability is leveraged using the LLM to generate subtasks from input instructions, which are then converted into word embeddings, denoted as $e_{w}$. The vertex features $F_{v}$ are the embeddings of each vertex’s textual description generated by the LLM. The graph adjacency matrix $A$ is a $2\times N$ matrix representing the connectivity between vertexes $x,y\in E(x,y)$. In LLMEG, the GNN encoder is implemented using two layers of graph convolution network (GCN) with the inputs as vertex features $F_{v}$ and graph adjacency matrix $A$. The encoder outputs graph embeddings $e_{g}$ of size $N\times F^{\prime}$, computed as:
$e_{g}=\sigma(WAF_{v})$ (1)
@@
where $N$ is the number of vertexes, $F^{\prime}$ is the dimension of the GCN layer, $W$ is a learned weighted matrix and $\sigma$ is the LeakyReLU activation function. The graph embedding is concatenated with the word embeddings to form the feature embedding $e_{f}=[\eta e_{g};e_{w}]$ as the input to the decoder, where $\eta$ is a trainable projection matrix that maps $e_{g}$ to the same latent space as $e_{w}$. The decoder consists of multi-layer perceptron (MLP) layers followed by a sigmoid activation function, producing the output:
$o=sigmoid(MLPs(e_{g}))$ (2)
@@
The output, with size $N$, indicates whether each vertex should be selected for the given task. To train the LLMEG encoder-decoder, the cross entropy loss $CE=-[vlog(\hat{v})+(1-v)log(1-\hat{v})]$ is used, where $v$ is the ground truth label indicating whether a vertex should be selected, and $\hat{v}$ is the predicted probability.
@@
The result $o$ from the LLMEG result is used to construct an undirected graph that identifies the APIs required to complete the task. To determine the sequence of API executions, classical path-searching algorithms, such as Depth First Search (DFS), can be applied. However, DFS lacks the ability to select the appropriate starting point due to its lack of semantic understanding. To address this limitation, the LLM is utilized again to transform the undirected graph into a directed graph $G^{*}$, incorporating semantic information from the subtasks. The system then executes the vertexes sequentially based on the order defined in the directed graph $G^{*}$.
III-D Ultrasound knowledge embedded and user intent understanding
@@
To adapt the model for understanding ultrasound tasks, a lightweight adaptation method, referred to as the LLaMA Adapter, is employed to efficiently fine-tune the original LLM into an instruction-following model specific to ultrasound-related tasks. The LLaMA Adapter achieves this by inserting prompted parameters into the top L layers of the N-layers LLaMA transformer model, enabling the model to acquire a new semantic understanding. In the USPilot framework, two adapters are trained: one based on ultrasound-related questions-answer pairs and the other tailored for ultrasound scanning tasks.
@@
Upon activating the LLaMA Adapter, the model experiences a reduced capability to handle general questions. As a comprehensive language model, retaining proficiency across a wide range of queries, including those outside the specific domain of robotic tasks is essential. However, the multi-adapter structure may compromise the model’s performance in general or medical question-answering scenarios while focusing on robotic tasks. A dynamic routing mechanism is integrated to address this limitation and restore versatility. The core component of this router is a three-layer MLP with a sigmoid activation gate. The complete architecture, including the dynamic router and adapters, is detailed in Fig4. The adaption prompt for the l-th layer of the LLaMA adapter can be defined as the concatenated instruction knowledge learned prompt $P_{l}$ and word tokens $T_{l}$ as
@@
, where the learned prompt $P_{l}$ acts as a prefix turning.The dynamic router manages and routes tasks effectively based on input. The learned adaption prompt can be illustrated as
$P_{l}=sigmoid(MLP(T_{n-l}))*P_{0}+sigmoid(MLP(T_{n-l}))*P_{1}$ (4)
@@
, where $P_{0}$ and $P_{1}$ represent prompt vectors from different adapters. The training dataset is divided strategically between general question-answering tasks and specific robotic task instructions. This division enables the dynamic router to adapt efficiently between these two domains, balancing specialized and general capabilities. The router mitigates overfitting by appropriately routing tasks, ensuring the model retains general applicability while specializing in robotic tasks. Furthermore, the architecture demonstrates the potential for routing multiple questions or tasks using multiclass activation functions such as softmax. In this case, the routing weights can be expressed as: $P=softmax(MLP(T_{n-l}))*[P_{0};P1;P_{2}...P_{n}]$.

### v2

@@ Abstract
In the era of Large Language Models (LLMs), embodied artificial intelligence presents transformative opportunities for robotic manipulation tasks. Ultrasound imaging, a widely used and cost-effective medical diagnostic procedure, faces challenges due to the global shortage of professional sonographers. To address this issue, we propose USPilot, an embodied robotic assistant ultrasound system powered by an LLM-based framework to enable autonomous ultrasound acquisition. USPilot is designed to function as a virtual sonographer, capable of responding to patients’ ultrasound-related queries and performing ultrasound scans based on user intent. By fine-tuning the LLM, USPilot demonstrates a deep understanding of ultrasound-specific questions and tasks. Furthermore, USPilot incorporates an LLM-enhanced Graph Neural Network (GNN) to manage ultrasound robotic APIs and serve as a task planner. Experimental results show that the LLM-enhanced GNN achieves unprecedented accuracy in task planning on public datasets with an accuracy of 78.64%, 60.8% and 59.6%. Additionally, the system demonstrates significant potential in autonomously understanding and executing ultrasound procedures with a successful demonstration of our physical setup with the robot. These advancements bring us closer to achieving autonomous and potentially unmanned robotic ultrasound systems, addressing critical resource gaps in medical imaging.
@@ I Introduction
Ultrasound is widely used in medical imaging. It provides real-time images of internal organs, blood vessels, and tissues without exposing patients to ionizing radiation. However, hospital sonographers face significant challenges due to the increasing demand for ultrasound services and the rapid pace of technological advancements. There are over 64,751 sonographers employed in the United States, but the number will need to increase by 27,600, according to a prediction by the Bureau of Labor Statistics[1]. In addition, sonographers commonly experience physical strain from their work. Repetitive motions often lead to joint discomfort, particularly in the wrists, upper back, and shoulders[2]. To mitigate this issue, it is imperative to consider the implementation of automation, particularly in light of the global shortage of ultrasound imaging professionals and the increasing demand for ultrasound-based diagnostic imaging.
The introduction of automated diagnostic and intraoperative ultrasound systems, which are either guided or assisted by robotic technology, represents a promising solution to the challenges previously outlined. There are currently several robotic assistant ultrasound scanning systems for specific tasks to enable more consistent digital images without sonographers,[3, 4] such as reinforcement learning[5, 6], imitation learning[7, 8], reward learning[4] and predefined tasks[9]. Specifically, Raina et al. propose a Bayesian optimization framework guided by an expert‑modeled “quality map” to direct probe movements[10]. Jiang et al. use a disentangled rewarding learning method to capture the “language of sonography” through self-supervised pairwise image comparisons[4]. They can achieve high accuracy, but these systems still require extensive domain knowledge or reward engineering, requiring doctors or engineers to select the specific functions individually. The human-robot interaction methods in these systems may also increase surgeons’ workload or distract them from their current tasks, which do not align with their operating habits and far away from future intelligent robotic sonographers[11].
@@
A subsequent surge in embodied artificial intelligence has been observed with the rise of large language models (LLMs) owing to their powerful associative capability and extensive prior knowledge. Some researchers have utilized its ability of understanding human text for task planning and reasoning, such as Octopus [12] and TAPA [13]. Other works go further on robot manipulation, such as [14, 15, 16] finetune the LLM on manipulating data together with pre-defined vision feedback. However, such low level robot control will easily face the illusion, which is unacceptable in medical scenarios.
In this paper, we present USPilot, a comprehensive framework designed for autonomous ultrasound acquisition. From the observation that when a human manipulates a device, they have an exact workflow in mind. The workflow is fixed due to the device being fixed, but it has various subpaths for different specific tasks. Hence, we defined the robotic ultrasound workflow as a graph and utilized the capability of LLMs to select a subgraph (i.e, a subpath as the human) based on the semantic information from the given task. Following this basic idea, the framework incorporates a semantic router powered by LLMs that interprets user queries related to ultrasound procedures and translates natural language for robotic ultrasound system operations. A key innovation of our framework is the implementation of an LLM-enhanced Graph Neural Network (GNN) for robotic tool selection and planning. By leveraging adapter-based ultrasound knowledge integration, USPilot demonstrates successful autonomous ultrasound image acquisition on a physical robotic platform, marking a significant step toward fully autonomous robotic ultrasound systems. This advancement suggests promising possibilities for unmanned medical imaging applications.
@@
II Methodology
II-A Problem formulation
@@
To build this task description graph $G$, note that a robotic control system naturally consists of low‑level functions, which can be exposed as related application programming interfaces (APIs). These executable APIs can form a graph where each vertex $v\in V$ corresponds to an API in the robotic system and each edge $(x,y)\in E$ indicates the dependency between two APIs. The robotic ultrasound manipulation can then be formulated as a task-planning problem on this fixed tool graph. However, unlike the structured graph, text instructions are in natural language, which is ambiguous and lies in another latent space. LLMs can perform as a perfect bridge as far as the API description text $t\in T$ is attached to each vertex to provide a linguistic description (eg. LiverSegmentation.Segments the liver from the ultrasound image.). Thus, the manipulation problem based on $I$ turns into a subgraph extraction on the predefined TaG graph $G=(V,E,T)$ according to textual information. In general speaking, the graph-based task planning aims to identify an optimal subgraph $G^{*}$ that can complete the given task instruction $I$.
@@ II-B Structure overview
The overview of USPilot is depicted in Fig. 2. The semantic router is a dynamic routing mechanism connected to the LLM transformer structure. It enables user intent recognition, which helps the system load the appropriate knowledge or graphs. Related subject knowledge is trained into the LLM through an adapter structure, allowing the model to better respond to user intent by adjusting the gating factor of the adapter layers.
@@
If the user intent involves executable tasks, the router switches to the Large Language Model Enhanced Graph (LLMEG) as Fig 3. The LLMEG comprises an LLM-enhanced Graph Neural Network (GNN) head and a subgraph generator. The LLM-enhanced GNN head selects APIs by sending LLM-generated word embeddings and GNN-generated graph embeddings to the decoder. The LLM-based subgraph generator defines the path for API execution. When the user intent involves a general medical question, the router switches the inference path to a pre-trained ultrasound-related question-answer adapter to fulfill the medical inquiry requirements. The semantic router, adapters, and LLMEG are all hot-plugging in our USPilot, which makes it suitable for separate training, upgrading, and adapting.
II-C Large language model enhanced graph
@@
LLMs possess a remarkable ability to interpret user requests and decompose complex commands into simpler subtasks. This task decomposition capability is leveraged using the LLM to generate subtasks from input instructions, which are then converted into word embeddings, denoted as $e_{w}$. To inherit the API description text into each vertex, inspired by LLaMA2 technical report[17] , it is suggested that the latent space of LLM is capable of handling the downstream tasks such like producing a scalar reward or hate speech classification, enriched by its pretrained prior knowledge. Accordingly, the vertex embedding $F_{v}$ of each vertex is extracted from LLMs’ last hidden layer to capture the textual features from inference conditioned on prompt and API description text. The graph adjacency matrix $A$ is an $N\times N$ matrix where nonzero entries indicate the presence of a connection, i.e., $A(x,y)\neq 0$ if $x,y\in E(x,y)$. The vertex embeddings and the adjacency matrix together form the graph structure fed into the GNN. In LLMEG, the GNN encoder is implemented using two layers of graph convolution network (GCN) with the inputs as vertex textual features $F_{v}$, and the graph adjacency matrix $A$. The encoder outputs graph embeddings $e_{g}$ of size $N\times F^{\prime}$, computed as:
$e_{g}=\phi(WAF_{v})$ (1)
@@
where $N$ is the number of vertexes, $F^{\prime}$ is the dimension of the GCN layer, $W$ is a learned weighted matrix and $\phi$ is the LeakyReLU activation function. The graph embedding is concatenated with the word embeddings to form the feature embedding $e_{f}=[\eta e_{g};e_{w}]$ as the input to the decoder, where $\eta$ is a trainable projection matrix that maps $e_{g}$ to the same latent space as $e_{w}$. The decoder $h(\cdot)$ consists of multi-layer perceptron (MLP) layers followed by a sigmoid activation function $\sigma$, producing the output:
$o=\sigma(h(e_{f}))$ (2)
@@
The output, with size $N$, indicates whether each vertex should be selected for the given task. To train the LLMEG encoder-decoder, the cross-entropy loss
$CE=-[vlog(\hat{v})+(1-v)log(1-\hat{v})]$ (3)
is used, where $v$ is the ground truth label indicating whether a vertex should be selected, and $\hat{v}$ is the predicted probability.
@@
The result $o$ from the LLMEG result is used to construct an undirected graph that identifies the APIs required to complete the task. To determine the sequence of API executions, classical path-searching algorithms, such as Depth First Search (DFS), can be applied. However, DFS cannot select the appropriate starting point due to its lack of semantic understanding. To address this limitation, the LLM is utilized again to transform the undirected graph into a directed graph $G^{*}$, incorporating semantic information from the subtasks. The system then executes the vertexes sequentially based on the order in the directed graph $G^{*}$.
II-D Ultrasound knowledge embedded and user intent understanding
@@
To adapt the model for understanding ultrasound tasks, a lightweight adaptation method, referred to as the LLaMA Adapter[18], is employed to efficiently fine-tune the original LLM into an instruction-following model specific to ultrasound-related tasks. The LLaMA Adapter achieves this by inserting prompted parameters into the top L layers of the N-layers LLaMA transformer model, enabling the model to acquire a new semantic understanding. In the USPilot framework, two adapters are trained: one based on ultrasound-related questions-answer pairs and the other tailored for ultrasound scanning tasks.
@@
Upon activating the LLaMA Adapter, the model experiences a reduced capability to handle general questions. As a comprehensive language model, retaining proficiency across a wide range of queries, including those outside the specific domain of robotic tasks is essential. However, the adapter structure may compromise the model’s performance in general or medical question-answering scenarios while focusing on robotic tasks. A dynamic routing mechanism $r(\cdot)$ is integrated to address this limitation and restore versatility. The core component of this router is a three-layer MLP with a sigmoid activation gate. The complete architecture, including the dynamic router and adapters, is detailed in Fig. 4. The adaptation prompt for the l-th layer of the LLaMA adapter can be defined as the concatenated instruction knowledge learned prompt $P_{l}$ and word tokens $T_{l}$ as
@@
, where the learned prompt $P_{l}$ acts as a prefix turning. The dynamic router manages and routes tasks effectively based on semantic input. The learned adaptation prompt can be illustrated as
$P_{l}=argmax(\sigma(r(T_{n-l})))*P_{0}+argmax(\sigma(r(T_{n-l})))*P_{1}$ (5)
@@
, where $P_{0}$ and $P_{1}$ represent prompt vectors from different adapters. The training dataset is divided strategically between general question-answering tasks and specific robotic task instructions. This division enables the dynamic router to adapt efficiently between these two domains, balancing specialized and general capabilities. The router mitigates overfitting by appropriately routing tasks, ensuring the model retains general applicability while specializing in robotic tasks. Furthermore, the architecture demonstrates the potential for routing multiple questions or tasks using multiclass activation functions such as softmax. In this case, the routing weights can be expressed as: $P=onehot(argmax(\sigma(r(T_{n-l}))))*[P_{0};P_{1};P_{2}...P_{n}]^{T}$, where $\sigma$ refers to the softmax activation function.
III Experiment
In this section, we present several experiments designed to assess the performance of the USPilot’s components. The LLMEG module was evaluated on a publicly available dataset compared to other methods. The semantic router and ultrasound knowledge adapter were also validated on a pre-collected dataset. Furthermore, the USPilot system was integrated with our robotic ultrasound scanning system, demonstrating its functionality in a physical environment.
III-A Datasets
To demonstrate that LLMEG module learns instructed API selection and task planning, three public datasets from TaskBench[19] were selected for evaluation, each with 500 test samples (1.5K in total). Multimedia: Contains 5.56K user-centric tasks such as file downloading and video editing. It is represented as a graph with 40 vertexes and 449 edges. HuggingFace: Covers various AI models covering 7.46K tasks across language, vision, audio, and video domains. Each task is treated as a tool for addressing specific problems, forming a graph with 23 vertexes and 225 edges. Dailylife APIs: Includes 4.32K tasks for daily services such as web search and shopping. It is represented as a graph with 40 vertexes and 1,560 edges.
The semantic router and ultrasound knowledge adapter were trained and evaluated on two self-created datasets as followings. We also created an anonymous repository to share the dataset in the URL: https://anonymous.4open.science/r/uspilot_data-0A0D/
Question Set: Contains 13.3K entries, which comprise 10.5K question-answer(QA) pairs and 2.8K robotic instructions. The 10.5K QA pairs combine 9K samples from the public Alpaca dataset[20] and 1.5K in-house self-collected ultrasound-related QA pairs. Each entry includes four keys: “instruction” contains user question, “input” contains additional context, “output” is the expected generation result of the LLM is not used in any training steps and “class” is the labeled user intent. The following is an example: {“instruction”: “Name 5 countries in the African continent.”, “input”: “”, “output”: “Nigeria, Egypt, South Africa, Ethiopia, and Ghana.”, “class”: “1”}.
Instruction Set: Contains 2.8K fully collected in-house instructions. Describes operational procedures of the ultrasound robotic system, mainly focusing on seven body parts: carotid vessel, thyroid, liver, kidney, spleen, femoral artery, and gallbladder. The dataset comprises instructions and expected subtasks, in which the instructions have the patient’s direct command and symptom description, such as “Scan the patient’s liver” and “The patient has jaundice; check him with ultrasound.”. The complexity varies from requiring to use 1 API to 9 APIs as shown in Fig.5 . Each entry includes three fields: “instruction” is the input for the LLM, “output” is the expected generation result of it and “API” is the ground truth of the robot API excitation, but it is not used in the fine-tuning stage. Here is an example of the instruction set: {“instruction”: “My left side feels tender to the touch.”, “input”: “”, “output”: “Step 1. Change the ultrasound probe to curvilinear probe. Step 2. Use camera to detect the spleen area on upper left part of abdomen. Step 3. Execute robotic scanning. Step 4. Segment spleen from real-time ultrasound image. Step 5. Publish report to gastroenterology department. Step 6. Done”, “API”: change_probe(curvilinear), detect_organ(spleen), execute_robot(), segment_organ(spleen), publish_report(gastroenterology)}.
Figure 5: Categorical distribution on organs and numbers of used APIs
III-B Comparison experiment for LLMEG
The LLMEG was trained on 5K training samples and 500 test samples from each of Multimedia, HuggingFace and Dailylife APIs. For optimization, we used the AdamW optimizer with a learning rate of $41e-$ 4, betas of $(0.9,0.95)$, a weight decay of 0.001, and a batch size of 64. We employed 3 metrics to evaluate the methods’ performance. Accuracy is the success rate of completing the given tasks. The vertex F1 score evaluates the labels assigned to individual vertexes, which can be indicated as
, where the $Precision_{vertex}$ is the number of correctly identified vertexes among all predicted vertexes. $Recall_{vertex}$ is the number of correctly identified vertexes out of the total actual vertexes, representing correct API calls. The edge F1 score measures the model’s ability to predict valid or expected API call sequences, using the same F1 equation
, where $Precision_{edge}$ and $Recall_{edge}$ follow the same definition as above, but applied on edges.